OpenStudy (anonymous):

An emission ray has a frequency of 5.10 × 10¹Ê Hz. Given that the speed of light is 2.998 × 10§ m/s, what is the wavelength of the ray?

OpenStudy (rushwr):

Hey @ataraiders Use this equation \[C = f \lambda \] Where C is Speed of light, f= frequency \[\lambda = wavelength \]

OpenStudy (rushwr):

SO substitute the values and subject wavelength to get the answer ! :)

OpenStudy (rushwr):

Did u get it ?

OpenStudy (anonymous):

its 15.2898 right

OpenStudy (rushwr):

Can u please show me the values again cuz i find it hard to calculate it with the numbers u have given in the question !

OpenStudy (rushwr):

powers of tens are not clear

OpenStudy (anonymous):

the first one is 10 to the 14th and the other is 10 to the 8th

OpenStudy (rushwr):

\[C= f \lambda \] \[\lambda = \frac{ C }{ f } = \frac{ 2.998 * 10^{8} }{ 5.1 * 10^{14} }= 5.878 * 10^{-7}\]

OpenStudy (rushwr):

I hope my values are correct

OpenStudy (anonymous):

it was right thank you

OpenStudy (anonymous):

would you be able to help me with one more question

OpenStudy (rushwr):

yeah open up a new question for that :)

OpenStudy (anonymous):

ok